# Log-Scale Quantization in Distributed First-Order Methods: Gradient-based Learning from Distributed Data

### v1

@@ IV-B Real Data-Set Example
First, we compare some existing algorithms in the literature with the proposed Algorithm 1 in the absence of quantization, i.e., we consider the linear case and set $h_{l}(z)=z$. This is only for the sake of fair comparison with the existing results. The following algorithms are used for comparison: GP [12], SGP [15, 16], S-ADDOPT [18], ADDOPT [17], and PushSAGA [24]. Fig. 5 presents the comparison results.
@@
Next, we compare training over exponential networks subject to uniform quantization versus log-scale quantization for different data quantization levels, see Fig. 6. As it can be observed, the logarithmic quantization results in a smaller optimality gap in the training outcome.
Figure 6: Optimality gap of uniform quantization (Top) versus log-scale quantization (Bottom) for distributed learning of MNIST dataset over exponential graphs.
The same comparison is performed over the geometric network and the result is shown in Fig. 7. As expected, logarithmic quantization results in a smaller optimality gap. For log-scale quantization in the simulation we could not use smaller values for $\rho$ as it returns the buffer overflow error (or stack overflow). This is due to limited memory of the computer program for this large-scale data-set as the calculation of $\exp\left(\rho\left[\dfrac{\log(|z|)}{\rho}\right]\right)$ in Eq. (4) returns stack overflow for small $\rho$. However, as illustrated in Fig. 4, the optimality gap can reach the order of $10^{-15}$ for log-scale quantization. It is worth noting that, as compared with the exponential network in Fig. 7, the optimality gap is generally larger over geometric networks as a non-structured topology.
Figure 7: Optimality gap of uniform quantization (Top) versus log-scale quantization (Bottom) for distributed learning of MNIST dataset over geometric graphs.

### v2

@@ Abstract
Note to Practitioners
Motivated by recent developments in cloud-computing and parallel processing along with low-cost CPUs and communication networks, this paper considers distributed/decentralized algorithms for machine learning and optimization. This work finds a wide range of practical applications in decentralized data mining, where the data set is distributed among a network of computing nodes. In the real-world scenarios, practical model nonlinearities such as data quantization must be addressed for information-exchange among the computing nodes. The primary emphasis of this work is to address log-scale quantization and compare it versus uniform quantization. Further, the impact of the structure of the data mining network is studied to reduce the optimality gap and improve the convergence rate. As contemporary distributed and networked data mining systems demand for more accurate algorithms with fast rate of convergence for real-time applications, our research highlights the intricate use of logarithmic quantization over structured networks. In particular, our results can be extended to other machine learning algorithms for more accurate and faster data mining solutions.
@@ IV-B Real Data-Set Example
First, we compare some existing algorithms in the literature with the proposed Algorithm 1. The following algorithms are used for comparison: GP [12], SGP [15, 16], S-ADDOPT [18], ADDOPT [17], and PushSAGA [24]. Fig. 5 presents the comparison results.
@@
Next, we compare training over exponential networks subject to uniform quantization for different data quantization levels, see Fig. 6. As it can be observed, the uniform quantization results in a larger optimality gap in the training outcome.
Figure 6: Optimality gap of uniform quantization for distributed learning of MNIST dataset over exponential graphs.
The same comparison is performed over the geometric network and the result is shown in Fig. 7. As expected, uniform quantization results in a larger optimality gap. It is worth noting from the figures that, the optimality gap is generally larger over geometric networks as a non-structured topology.
Figure 7: Optimality gap of uniform quantization for distributed learning of MNIST dataset over geometric graphs.